OpenStudy (anonymous):

List all the subsets of {3, 4, 7}. For this, would one of the subsets be {} (blank)? And would the others be the numbers alone:{3}, {4}, {7}. And then the pairs:{3,4}, {3,7}, {4,7}. Right? Just want to make sure I'm doing it right! Thanks!

OpenStudy (anonymous):

No of subsets = 2^3 = 8

OpenStudy (anonymous):

\[\phi \]

OpenStudy (anonymous):

What? Sorry.. I just don't get what your saying..

OpenStudy (anonymous):

No of subsets =2 ^ no of elements in the given set

OpenStudy (anonymous):

{{3}, {4}, {7} , {3,4} , {3,7}, {4,7} , {3,4,7} , phi}

OpenStudy (anonymous):

What is phi? & Thanks!

OpenStudy (anonymous):

You're correct. You got the single element sets and the sets with pairs. Yes, {} is a subset also =) It's the empty set, meaning it has 0 elements, and it's often written as \(\large \emptyset\). The empty set is a subset of every other set. The only one you missed is the set itself {3, 4, 7}. Technically, it is also a subset. A set is always considered to be a subset of itself.

OpenStudy (anonymous):

Oh okay, thank you :)

OpenStudy (anonymous):

My pleasure =D